### v1

# Towards Accurate and Efficient 3D Object Detection for Autonomous Driving: A Mixture of Experts Computing System on Edge

Abstract
This paper presents Edge-based Mixture of Experts (MoE) Collaborative Computing (EMC2), an optimal computing system designed for autonomous vehicles (AVs) that simultaneously achieves low-latency and high-accuracy 3D object detection. Unlike conventional approaches, EMC2 incorporates a scenario-aware MoE architecture specifically optimized for edge platforms. By effectively fusing LiDAR and camera data, the system leverages the complementary strengths of sparse 3D point clouds and dense 2D images to generate robust multimodal representations. To enable this, EMC2 employs an adaptive multimodal data bridge that performs multi-scale preprocessing on sensor inputs, followed by a scenario-aware routing mechanism that dynamically dispatches features to dedicated expert models based on object visibility and distance. In addition, EMC2 integrates joint hardware-software optimizations, including hardware resource utilization optimization and computational graph simplification, to ensure efficient and real-time inference on resource-constrained edge devices. Experiments on open-source benchmarks clearly show the EMC2 advancements as a end-to-end system. On the KITTI dataset, it achieves an average accuracy improvement of 3.58% and a 159.06% inference speedup compared to 15 baseline methods on Jetson platforms, with similar performance gains on the nuScenes dataset, highlighting its capability to advance reliable, real-time 3D object detection tasks for AVs.
1 Introduction
Traffic safety is a fundamental concern for both human drivers and Autonomous Driving Systems (ADS). Within an ADS, the perception module plays a pivotal role by serving as the ”eyes” of the autonomous vehicle (AV), sensing the surrounding environment and delivering critical information to downstream modules such as motion planning, control and safety guarantees. Existing conceptual safety of an AV perception system is generally defined by two core dimensions: accuracy and efficiency. Accuracy ensures reliable object detection and tracking capabilities, enabling a credible understanding of the environment. Efficiency, characterized by low perception delay, is equally essential, as real-time high-level planning and control rely on fast information processing. Perception delay, defined as the sum of data acquisition, computation, and algorithm processing time, directly affects an ADS’s responsiveness to dynamic environments. Research indicates that an ADS must process sensory inputs and issue feedback control commands within milliseconds to guarantee prompt and safe vehicle operations [1]. Detection precision and perception delay together define a critical safety boundary: a precise but sluggish system cannot react in time, while a fast but inaccurate system may lead to unsafe decisions.
As illustrated in Fig. 1, achieving both high accuracy and low latency remains a fundamental challenge. Existing approaches often face inherent trade-offs between these objectives due to limited computational resources, insufficient adaptability of perception systems, and the inherent complexity of dynamic driving scenarios. Urban traffic environments require low-latency and accurate sensing capabilities to handle complex, densely interactive surroundings, while highway scenarios demand reliable detection of distant objects and prompt responses to unexpected events. Moreover, even for the same object, recognition priorities may vary dynamically within a scene depending on distance and scene complexity, posing notable challenges for single-modality models in balancing efficiency and accuracy [2]. Multimodal models, such as large-scale CNN and Transformer-based architectures that integrate LiDAR, radar, and camera data [3], leverage the complementary advantages of diverse sensors to enhance perception accuracy. Nevertheless, these approaches [4] often impose substantial hardware overhead and exhibit limited adaptability, particularly when deployed on resource-constrained edge platforms [5]. Even with the assistance of low-power machine learning compilation frameworks [6], such systems may still execute redundant computations under straightforward scenarios, introducing additional delays that undermine timely safety-critical decision-making.
Nevertheless, multimodal approaches provide a natural foundation for addressing the accuracy challenges faced by ADS. Rather than relying on a single integrated model designed to operate uniformly across diverse driving scenarios, the MoE framework [7] dynamically activates specialized sub-models, or experts, based on input characteristics and prior knowledge. MoE has demonstrated considerable effectiveness in large-scale applications, such as Large Language Models (LLMs) for natural language processing [8, 9], where it routes different tokens to appropriate experts, reducing redundant computations without sacrificing efficiency [10]. As model complexity grows, MoE enables models to scale their capacity while keeping inference hardware costs manageable. Deploying MoE in ADS, however, poses notable challenges [11], including strict real-time constraints, the need for multimodal data processing, and seamless end-to-end autonomous feedback control. Successful deployment requires efficient data preprocessing, adaptive modality selection, and robust expert routing mechanisms.
By leveraging diverse sensor data and a scenario-adaptive MoE dispatcher, we dynamically assign expert sub-models based on scenario data and prior statistical knowledge, providing an efficient, energy-efficient, and safe solution for ADS. To this end, we introduce an Edge-based MoE Collaborative Computing () framework – an algorithm–hardware–software co-design solution that not only improves accuracy through complex multimodal experts but also effectively reduces latency by utilizing an adaptive MoE dispatcher and lightweight single-modal experts. Our main contributions are summarized as follows:
A Multimodal 3D Object Detection Solution with MoE. The proposed system achieves both high detection accuracy and inference efficiency, outperforming existing one-size-fits-all baseline models. Designed for real-world ADS applications, is optimized for deployment on resource-constrained platforms, making it more suitable for edge computing environments than generic alternatives.
Robust and Efficient Multimodal Perception through Expert Collaboration and System-Level Optimization. The proposed framework incorporates scenario-specific experts and employs an adaptive expert selection strategy to meet heterogeneous perception demands. To mitigate the training instability commonly observed in MoE-based systems, we design a hierarchical optimization scheme that combines expert-level and global backpropagation for effective collaborative learning. Furthermore, a distribution-aware resampling technique enhances robustness under real-world data imbalance. In addition to these designs, a tailored multiscale pooling module improves the efficiency of LiDAR–camera fusion, ensuring both efficient and accurate multimodal perception.
System-Level Optimization for Edge Inference. The original PyTorch-based MoE inference framework [12] suffers from inefficient management of memory and computational resources, leading to high latency on edge devices. To overcome these challenges, we propose an algorithm–hardware–software co-optimization strategy within , which incorporates efficient memory management techniques to improve L1/L2 cache hit rates and applies computational graph fusion to eliminate redundant data loading across layers. These combined optimizations notably improve inference efficiency.
SOTA Accuracy and Efficiency Combined. Experiments on the KITTI dataset demonstrate that clearly improves detection accuracy for challenging (hard-level) objects, achieving 3.28% higher accuracy for pedestrians and 6.03% for bicycles compared to baseline models. On the nuScenes dataset, further achieves a 3.9% improvement in mAP and a 1.8% boost in NDS over existing SOTA solutions. Moreover, delivers 159.06% faster inference on a Jetson platform on the KITTI benchmark, substantially enhancing real-time processing capabilities without sacrificing detection performance, and making it well-suited for edge deployment.
2 Related Work
LiDAR-only Method. The active 3D sensing of LiDAR makes it dominant in ADS 3D detection [13, 14, 15]. LiDAR-based methods process point clouds via voxelization [16], pillarization [15], or direct raw input [17]. Anchor-based heads, used for region proposal [16], and center-based representations[18] are widely adopted. These methods achieve high accuracy on KITTI, as demonstrated by 3DSSD [19], PartA2 [14], and Voxel R-CNN [3]. However, weak LiDAR reflections for distant or occluded objects degrade detection performance [20, 21].
Multimodal Fusion Method. Combining LiDAR and camera RGB input enhances detection performance, especially for sparse LiDAR objects [22]. Fusion approaches include projection-based methods, which align image features with point clouds in 3D space [23], and model-based techniques that incorporate cross-attention [24] or unified feature spaces [25]. Recent multimodal models, such as PPF-Det [26] and RoboFusion [27], achieve top accuracy on KITTI benchmarks. However, these models often trade off accuracy for speed or vice versa.
Edge-based Compilation. ML compilation platforms are essential for improving the inference efficiency of deep learning models across diverse hardware architectures [28, 29]. TensorRT [30] is an NVIDIA inference optimizer that accelerates ML models on GPUs through layer fusion, precision calibration, and kernel auto-tuning. XLA [31] is a domain-specific compiler for TensorFlow enhancing efficiency by fusing operations and optimizing computation graphs. Open Neural Network Exchange (ONNX) Runtime [32] provides an open standard for model interoperability, enabling efficient execution across CPUs, GPUs, and accelerators. TVM [6] is an open-source deep learning compiler that automates model tuning and tensor scheduling for efficient deployment across CPUs and GPUs.
3 EMC2 Computing Methodology
We introduce , a multimodal MoE framework that minimizes latency and ensures reliable accuracy. Fig. 2 shows the overview, consists of: 1) Adaptive Multimodal Data Bridge (AMDB) to preprocess multimodal input (Sec. 3.1); 2) Scenario-Adaptive Dispatcher to adaptively dispatch a suitable expert for each scenario (Sec. 3.2); and 3) Three Scenario-Optimized Experts to decode AMDB-output features into final results (Sec. 3.3). We also design an effective training strategy to address the long-tail effect in the training process (Sec. 3.4).In addition, we deploy on edge devices with self-designed 3D sparse convolution and Multiscale Pooling.
3.1 Adaptive Multimodal Data Bridge
Given the multimodal nature of ’s sensor inputs, a module providing suitable preprocessed data for different experts is required. Thus, we propose AMDB, a preprocessing unit to efficiently process the required input data for each expert. The workflow of AMDB begins with a UNet [33] and a sparse CNN to extract LiDAR features, followed by fully connected layers that generate proposal regions along with their confidence scores. These results serve as inputs to Latency-Prioritized Expert and Versatile Efficiency Expert. The extraction of image features depends on the Scenario-Adaptive Dispatcher. If further information is required for reliable detection, AMDB will extract image features by ResNet. Based on depth information, it projects relevant image pixels into 3D space, applying Multiscale Pooling to reduce computational overhead, and fuses them with LiDAR voxel features to produce a multimodal output. These results are inputs for Accuracy-Prioritized Experts. The Multiscale Pooling and multimodal fusion processes are demonstrated in Fig. 5.
3.2 Scenario-Adaptive Dispatcher
The Scenario-Adaptive Dispatcher (SAD) dynamically dispatches each scenario to the most suitable expert based on traffic conditions and latency-accuracy requirements. Our expert selection algorithm considers real-world traffic safety needs: specifically, how far away objects are and how clearly they can be perceived.
So, the switch among experts is determined by two key parameters: object distance and clarity, which is inferred from the confidence of proposal regions provided by the AMDB. We classify traffic scenarios into three categories: 1) Close and Distinct Cases. All objects are clearly visible and no distant objects are present. This is characterized by all proposal regions and having confidence . In such scenarios, as near-field objects provide abundant LiDAR information, even when compressed into bird’s eye view (BEV), a 2D CNN can effectively capture key patterns such as object surfaces and heights [15], ensuring reliable detection accuracy with minimal latency (Sec. 4.5). These cases will be routed to the Latency-Prioritized Expert. 2) Mixed Visibility Cases. Some objects are either distant but clearly visible or near but unclear. Mathematically, this corresponds to cases where i) some proposal regions with confidence , or ii) the distance of some objects with confidence . Here, objects have insufficient voxel data, requiring more complex 3D convolution to extract meaningful patterns, and they will be dispatched to the Versatile Efficiency Expert. 3) Distant and Uncertain Cases. Some objects are both distant and unclear. This occurs when some proposal regions with confidence . In these challenging scenarios, significant voxel information is missing due to long distances, steep reflection angles, or occlusions, necessitating the integration of image data to compensate for missing LiDAR details. These cases will be processed by the Accuracy-Prioritized Expert.
3.3 Scenario-Optimized Experts
Latency-Prioritized Expert (LPE). LPE utilizes 2D-CNN for object detection based on BEV projection. It first projects proposal regions in 3D voxel space, generated by the AMDB, into a BEV representation. Next, it extracts features and performs object localization and classification using a 2D CNN before mapping the detection results back into 3D space. Compared to sparse 3D convolutions, 2D convolutions offer significantly lower computational costs in outdoor ADS scenarios [15]. By leveraging this efficiency, LPE achieves high inference efficiency while maintaining reliable detection accuracy.
Versatile Efficiency Expert (VEE). VEE integrates a 3D CNN to process proposed regions and confidence scores from the AMDB. After sparse convolution, the extracted voxel features are passed through a trained decoder that generates per-voxel classifications and bounding boxes as the result. Compared to LPE, the use of 3D convolutions allows more accurate spatial structure preservation, making it more robust for scenarios with insufficient LiDAR signal.
Accuracy-Prioritized Expert (APE). APE extends VEE by incorporating the multimodal signal. It employs a 3D CNN and decoder with the same structure but different parameters to process multimodal proposal regions (from AMDB) and fuses image data to recover missing LiDAR information. Image features compressed via Multiscale Pooling (Sec.3.5) are projected into 3D space and fused with corresponding voxels of LiDAR features, ensuring more robust understanding for occluded and distant objects while keeping computation feasible. The Multiscale Pooling and multimodal fusion are detailed in Fig. 5.
3.4 Collaborative Training of Multimodal Experts
Hierarchical Training Strategy: Multimodal Supervision with Triple Back-Propagation. During training, the performance of experts is highly dependent on the proposal regions produced by the AMDB, yet these proposals tend to be unreliable in the early training phase, leading to an unstable expert learning process and unstable SAD decisions. To address this, we adopt a hierarchical supervision strategy with three distinct back-propagation pathways, as illustrated in Fig. 4: 1) supervision of the LiDAR branch outputs from the AMDB; 2) supervision of the image branch outputs from the AMDB; and 3) supervision of expert predictions and associated confidence scores. To further improve training stability, the LiDAR and image branches of the AMDB are pre-trained separately before joint optimization. Each back-propagation route is supervised by the following loss function, which collectively forms Multimodal Loss:
where uses cross-entropy loss and uses smooth- loss. The Hierarchical Training not only stabilizes expert learning but also enhances the model’s adaptability to various data granularities. In turn, even under compressed data representations, it enables the effective use of our Multiscale Pooling for memory efficiency, as detailed in Sec. 3.5.
Addressing Long-tail Effects in MoE Training. The MoE framework often suffers from long-tail effects, where certain experts receive insufficient training data, leading to imbalances in sub-datasets, and therefore affecting performance [34]. To address this, we propose a threefold training strategy: 1) Data subset division. We partition the dataset into sub-datasets based on target and auxiliary samples. Each expert is dispatched a sub-dataset () containing target samples , which are expected to be routed to the corresponding expert, and auxiliary samples , which are not. We ensure that for any . The inclusion of auxiliary samples, which are not intended for a given expert, helps prevent overfitting and enhances generalization. 2) Balanced sampling strategy. We adjust the selection probability for the sub-datasets which have fewer samples to ensure balanced training:
where represents the number of samples in sub-datasets . This ensures that experts dispatched to less frequent scenarios receive sufficient training iterations. 3) Adaptive optimizer. We introduce an optimizer that adapts the learning rate based on the proportion of target samples in a batch. Denoting as the base learning rate, the adaptive learning rate applied to an expert with a given batch is computed as:
where is the batch size, is the proportion of target samples within current mini-batch, represents target samples in the batch. By regulating the gradient update rate based on the batch composition, this adaptive optimizer encourages each expert to prioritize its designated target samples while preventing excessive updates from auxiliary samples.
3.5 Algorithm Edge-adaptivity Optimization
Edge devices face challenges such as incomplete function libraries and limited cache. We adopt the open-source ONNX Runtime for compiling on edge, which provides modules that help address these challenges and facilitate edge-oriented optimization. As 3D convolutions account for the majority of computation during inference, we customize 3D sparse convolution and Multiscale Pooling to improve the algorithmic efficiency of .
3D Sparse Convolution. Since the 3D sparse convolution library is incomplete on most edge platforms, we develop a customized 3D sparse convolution library with parallel threads execution under ONNX Runtime. Let denote spatial resolution, the number of non-empty voxels, and the per-voxel computation, including channels and kernel size. By computing nonzero voxels only, 3D sparse convolution reduces complexity from for dense convolutions to , where , cutting redundant operations by 65–80% [35].
Multiscale Pooling. To alleviate cache limitations, Multiscale 3D Sparse Pooling adaptively pools image features based on available memory before fusing them with LiDAR features (Fig.5). The pooling size is user-configurable, allowing flexibility for maximizing memory utilization. When combined with Hierarchical Training (Sec.3.4), Multiscale Pooling effectively reduces resource consumption while maintaining model reliability (Sec. 4.5).
3.6 EMC2 Computing System Optimization
Building upon the algorithm-level optimizations introduced in Sec. 3.5, we further enhance the ONNX execution efficiency through system-level improvements, specifically targeting memory management and computation graph optimization.
Memory Optimization. To further improve memory efficiency during inference, we implement three system-level techniques. 1) Overlapping communication and computation [36]. Matrix data is partitioned into multiple segments, which are processed and transferred in parallel across processors. This overlapping of data movement and computation effectively reduces global memory access and improves cache utilization, alleviating potential memory bottlenecks. 2) Thread scheduling. We implement a staged thread management mechanism, where active threads are periodically terminated and re-initialized (Fig. 6). This approach limits unnecessary memory retention between stages, reducing system memory footprint. 3) Prefix-sum acceleration for sparse convolution. To speed up the index search for nonzero elements in sparse tensors, we apply parallel prefix-sum operations, reducing the time complexity from to by leveraging GPU parallelism, which boosts sparse convolution efficiency, and denotes the number of nonzero tensor elements.
Computational Graph Optimization. ONNX partitions the computation graph into multiple segments, dispatching each to the most suitable hardware block. Our optimization includes three stages: 1) Model pruning. Redundant parameters and operations are removed to simplify the computational graph structure. 2) Model quantization. Low-precision computations replace floating-point operations to improve hardware resource efficiency. 3) Computational graph fusion. Consecutive operators are merged to reduce memory access overhead and improve execution efficiency, which is particularly beneficial for memory-bound hardware such as CPUs, and also improves kernel launch efficiency on GPUs.
4 Experiment and Result
4.1 Experiment Settings and Datasets
Experiment Platform. We conduct experiments on the Jetson AGX Orin edge platform, which features limited computational resources, an NVIDIA Ampere GPU, 4MB shared L2 cache, and a hierarchical memory structure. Due to the absence of advanced sparse convolution libraries, matrix operations that exceed the L2 cache size result in notable memory exchange overhead, severely affecting computational efficiency. In addition, for comparison, we also run inference on an NVIDIA A4000 GPU.
Dataset and Sample Difficulties. We evaluate the 3D object detection performance of using two public benchmarks: KITTI [37] and nuScenes [38]. The KITTI dataset contains training samples and validation samples. According to the official KITTI difficulty levels, of the objects are labeled as “easy,” as “moderate,” and as “hard.” The nuScenes dataset consists of training scenes and test scenes, with object categories and 3D bounding box annotations for 3D detection.
Evaluation Metrics. For the KITTI dataset, detection accuracy is evaluated using the R40 3D Average Precision (AP) metric, with Intersection over Union (IoU) thresholds of for cars and for pedestrians and cyclists. The nuScenes dataset adopts the NDS metric, which combines mAP with a set of true positive error metrics. The mAP is computed based on center distance matching rather than IoU, with distance thresholds ranging from m to m depending on the object category. The true positive error metrics include mean Average Translation (mATE), Scale (mASE), Orientation (mAOE), Velocity (mAVE), and Attribute (mAAE) Errors. In addition to detection accuracy, inference latency is measured on the Nvidia Jetson AGX Orin to assess real-time deployment capability.
4.2 Implementation Details
Hierarchical Training Details. During training, we observed that allowing all experts to propagate gradients into the AMDB leads to instability. This is because the training data for different experts partially overlap, causing AMDB to be corrected multiple times for the same errors [7]. To address this, we first pre-train AMDB independently for epochs before integrating the experts. We then enable joint training with all three experts while restricting gradient propagation to AMDB solely from the APE, as APE is the most comprehensive expert that incorporates multimodality.
Data Subset Division Details. Prior to the data subset division, we conduct a statistical analysis of the confidence distribution of proposal regions generated by the aforementioned pre-trained AMDB model. The results of the two-sample Kolmogorov-Smirnov test [39] reveal a statistically notable difference in the confidence distributions of objects located within and beyond a certain distance . Specifically, the distance threshold is set to meters for the KITTI dataset and meters for the nuScenes dataset, based on empirical observations from each dataset. Accordingly, we define the target subsets for training based on the specialization of each expert as follows: 1) For the LPE, the target subset consists of scenes where all objects are within meters and labeled as “easy” or “moderate.” 2) For the VEE, the target subset includes scenes where at least one object is beyond meters but remains classified as “easy” or “moderate,” or where objects within meters are labeled as “hard.” 3) For the APE, the target subset contains scenes in which at least one object is simultaneously beyond meters and labeled as “hard,” representing the most challenging cases requiring multimodal fusion. For each expert, the auxiliary subset is uniformly sampled from the target subsets of the other experts. The number of auxiliary samples matches the number of target samples in each subset.
Adaptive Optimizer Details. We observed that when batch size is too small, the learning rate fluctuates abruptly due to the binary nature of target sample presence, leading to unstable training. Conversely, when is too large, the learning rate remains nearly constant, reducing its effect in prioritizing target samples. Through empirical evaluation, we choose a moderate batch size of to ensure the intended effect of Adaptive Optimizer while maintaining smooth learning rate adjustments.
| Method | Pedestrian 3D AP (R40) | Car 3D AP (R40) | Cyclist 3D AP (R40) | Latency (ms) | |||||||
| Easy | Mod. | Hard | Easy | Mod. | Hard | Easy | Mod. | Hard | A4000 | Jetson | |
| LiDAR-only Models | |||||||||||
| M3DETR [40] | 69.69 | 66.04 | 61.61 | 93.96 | 86.28 | 84.32 | 87.88 | 72.47 | 70.63 | 338 | N/A |
| PV-RCNN [41] | N/A | N/A | N/A | 91.18 | 87.65 | 83.14 | N/A | N/A | N/A | 45 | 1787 |
| Voxel-RCNN [3] | N/A | N/A | N/A | 92.23 | 85.04 | 82.50 | N/A | N/A | N/A | 70 | N/A |
| GLENet-VR [42] | N/A | N/A | N/A | 93.51 | 86.22 | 83.72 | N/A | N/A | N/A | 165 | N/A |
| PartA2-Anchor [43] | 66.87 | 59.68 | 54.60 | 92.15 | 82.92 | 82.10 | 90.34 | 70.05 | 66.89 | 95 | N/A |
| PartA2-Free [43] | 72.31 | 66.36 | 60.06 | 91.66 | 80.28 | 78.08 | 91.88 | 75.33 | 70.67 | 124 | 1064 |
| PointPillar [15] | 57.29 | 51.41 | 46.87 | 87.75 | 78.40 | 75.19 | 81.57 | 62.93 | 58.97 | 41 | 972 |
| Second [35] | 55.94 | 51.14 | 46.16 | 90.55 | 81.60 | 78.60 | 82.96 | 66.73 | 62.78 | 45 | 1322 |
| Second-IoU [35] | 61.10 | 54.66 | 49.50 | 91.53 | 82.36 | 79.62 | 90.73 | 71.23 | 66.26 | 58 | N/A |
| 3ONet [44] | 72.55 | 65.21 | 60.22 | 94.24 | 87.32 | 84.17 | 92.47 | 75.11 | 71.18 | 100 | N/A |
| Multimodal Models | |||||||||||
| Voxel-RCNN [3] | N/A | N/A | N/A | 92.08 | 85.90 | 83.36 | N/A | N/A | N/A | 331 | 965 |
| LoGoNet [45] | 70.02 | 63.72 | 59.46 | 92.04 | 85.04 | 84.31 | 91.74 | 75.35 | 72.42 | 100 | N/A |
| VFF [46] | 73.26 | 65.11 | 60.03 | 92.24 | 85.51 | 82.92 | 91.74 | 75.35 | 69.84 | 192 | N/A |
| CAT-DET [47] | 74.08 | 66.35 | 58.92 | 90.12 | 81.46 | 79.15 | 87.64 | 72.82 | 68.20 | 154 | N/A |
| TED [48] | 74.73 | 69.07 | 63.63 | 92.25 | 88.94 | 86.73 | 95.20 | 76.17 | 71.59 | 162 | N/A |
| (Ours) | 74.92 | 69.92 | 66.81 | 94.00 | 89.35 | 88.15 | 95.06 | 79.87 | 77.62 | 160 | 372.5 |
4.3 Comparison with Existing SOTA Methods
As shown in Tab. 1 and Tab. 2, achieves SOTA performance in terms of both detection accuracy and inference latency on the KITTI and nuScenes validation sets. Furthermore, demonstrates superior performance on the Jetson AGX Orin platform, outperforming existing models across various difficulty levels.
Results on KITTI. Pedestrians and cyclists, as vulnerable road users, face higher risks in traffic [49]. However, most SOTA methods underperform in detecting these categories effectively. In Tab. 1, improves pedestrian and cyclist detection accuracy by and , respectively, on the hard setting, outperforming recent methods such as TED and LoGoNet. For car detection, achieves a robust accuracy of under the hard setting, surpassing previous multimodal approaches by up to . Importantly, demonstrates practical deployability with an average latency of ms on the Jetson AGX Orin for KITTI, representing a to speedup over existing methods that require ms to ms.
Results on NuScenes. In Tab. 2, achieves improvements of and in mAP and NDS, respectively, outperforming the latest methods FocalFormer [50] and BEVFusion [51]. Meanwhile, delivers an average inference latency of ms on the Jetson AGX Orin for nuScenes, representing a high inference speed for real-time edge deployment.
4.4 Analysis and Discussion
Benefits of Mixture-of-Experts. The balance between accuracy and efficiency demonstrated in Sec. 4.3 is achieved by dynamically selecting experts based on scenario characteristics rather than adopting a one-size-fits-all approach. This is further supported by comprehensive algorithmic and system-level optimizations tailored for edge devices with limited computational resources, such as Jetson.
Empirical Dispatcher. The dispatcher of employs a lightweight and interpretable decision-making mechanism based on fixed, empirically determined statistics, specifically distance and confidence thresholds, dynamically dispatches each scene to the most appropriate expert. Due to dataset format variations, the dispatch protocol is designed individually for each dataset. To validate this approach, we analyzed expert activations during inference on the nuScenes validation subset, as shown in Fig. 7. LPE achieves the best performance in scenes with close-range objects and high confidence scores, APE excels in distant, low-confidence scenarios, and VEE outperforms the others in all remaining cases. These observations on expert selection support the effectiveness of using empirical statistics for our dispatcher’s routing strategy.
Optimized Parameter Size. At the algorithmic level, the full contains 226M parameters in total. However, with the MoE framework, only a subset of parameters is used during each inference – averaging 87M on nuScenes and 146M on KITTI. This is substantially lower than the original and other SOTA models. At the system level, our optimization focuses not on reducing parameters, but on maximizing hardware efficiency by accelerating computation and minimizing memory access overhead.
Hardware Flexibility. The is implemented within the ONNX format, an open format supported by compilers that enables model compilation across diverse hardware platforms. As a result, is not limited to Jetson and can be deployed on other edge platforms.
4.5 Ablation Study
| Method | AP | Latency (ms) | ||
| Ped. | Car | Cyclist | ||
| Ablation on LPE | ||||
| LPE | 66.32 | 89.04 | 88.54 | 219 |
| VEE | 74.49 | 92.23 | 91.69 | 468 |
| APE | 74.34 | 93.21 | 92.41 | 526 |
| Ablation on VEE | ||||
| W/ VEE | 70.55 | 90.50 | 84.07 | 372.5 |
| W/O VEE | 68.26 | 86.40 | 79.17 | 512.7 |
| Ablation on APE | ||||
| VEE | 67.24 | 87.69 | 81.85 | 468 |
| APE | 70.55 | 90.50 | 84.18 | 526 |
| Ablation on Hierarchical Training and Multiscale Pooling | ||||
| HT ✓, MP ✓ | 70.55 | 90.50 | 84.07 | 526 |
| HT ✓, MP ✗ | 70.55 | 90.50 | 84.18 | 1320 |
| HT ✗, MP ✓ | 63.11 | 88.68 | 75.98 | 529 |
| HT ✗, MP ✗ | 66.34 | 89.10 | 80.83 | 1367 |
Ablation on the Role of LPE. We further investigate the efficiency advantages of LPE compared to VEE and APE under close-range and clearly visible cases. Specifically, we assess three model variants: 1) LPE, 2) VEE, and 3) APE, all evaluated on the KITTI validation set corresponding to LPE’s target scenarios. As summarized in the first sub-table of Tab. 3, all three maintain comparable detection performance, while LPE notably reduces inference time on the Jetson platform, highlighting its effectiveness for low-latency deployment in suitable conditions.
Ablation on the Impact of VEE. We compare the performance of with and without VEE using the KITTI validation subset corresponding to VEE scenarios. The results, presented in the second sub-table of Tab. 3, demonstrate that incorporating the VEE expert leads to clear improvements in both detection accuracy and inference efficiency compared to the baseline.
Ablation on the Effect of APE. To assess the benefits of incorporating multimodal information, we conduct a comparative evaluation of VEE and APE on the full KITTI validation set. As reported in the third sub-table of Tab. 3, integrating the Image branch from the AMDB leads to a notable improvement in detection accuracy, with only a slight increase in inference time on the Jetson AGX Orin. This observation is further supported by the detection score distribution illustrated in Fig. 8.
Ablation on the Impact of Hierarchical Training (HT) and Multiscale Pooling (MP). As discussed in Sec. 3.4 and Sec. 3.5, HT directly supervises the two branches of AMDB during training and facilitates the application of MP, which enhances inference efficiency. We evaluate the combined impact of these strategies through four settings: 1) training with HT and inference with MP; 2) training with HT only; 3) inference with MP only; and 4) neither strategy used. All settings are trained for the same number of epochs and evaluated on the full KITTI validation set. The results, shown in the third sub-table of Tab. 3, indicate that although accuracy appears slightly higher when MP is not applied, combining it with HT achieves a reduction in inference time on Jetson, from ms to ms, without compromising accuracy. This is attributed to the complementary strengths of HT, which effectively learns representations at multiple levels of granularity, and MP, which leverages data compression to optimize memory utilization on resource-constrained devices. However, solely applying MP significantly reduces accuracy from to . Moreover, even employing HT alone substantially enhances accuracy from to under the same training epoch.
5 Conclusion
We introduced , a multimodal 3D object detection solution for ADS, built upon a customized MoE-based architecture. This design enables to dynamically adapt to varying traffic conditions while maintaining a practical trade-off between detection accuracy and computational efficiency. Extensive experiments on the KITTI and nuScenes datasets demonstrate that achieves state-of-the-art performance, with notable improvements under challenging conditions. Moreover, the framework delivers a inference speedup on the Jetson platform, making it highly suitable for real-time edge deployment. In the future, we will develop a fully adaptive expert selection mechanism to further improve system adaptability and edge deployment efficiency.